OpenStudy (anonymous):

Find the distance between (-2, 8) and (11, 2). Round to the nearest hundredth

OpenStudy (anonymous):

use distance formula

OpenStudy (anonymous):

what is the distance formula?

OpenStudy (anonymous):

\[\sqrt{\left( \left( x _{2}-x _{1}^{2} \right)+\left( y _{2}-y _{1}^{2} \right) \right)}\]

OpenStudy (anonymous):

I accidentally placed the ^2 inside the parenthesis when it's supposed to be outside both parenthesis

OpenStudy (anonymous):